Session: Feb 2005 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Roda at Mon Mar 7 11:01:02 2005) Success :

Torun (K. Borkowski at Mon Mar 7 12:00:42 2005) Minor failures :
BBC3 unlocked (same problem as in GD018). No other known problems.

Jodrell_Bank (Alastair Gunn at Mon Mar 7 12:35:50 2005) Minor failures : LATE
MkII was used. Lovell telescope drive still faulty, so schedule was run using the Mk2 telescope instead. Error at start of experiment turned out to be antenna control computer needing reboot. Several FS restarts to try and clear this problem. Schedule was finally on source with correct LOs for 23:20 UT scan. Phase CAL intermittently on and off during this experiment. No other known problems.

Westerbork (Tony Foley at Mon Mar 7 15:27:27 2005) Minor failures : WIND HIGH PHAS
Wb used in phased array mode. Recabled at the start for Mk4. At the end of this observation most of the snow in the dishes was melting fast. It looks like the phasing up was bad in IVC band 2 (=BBCs 5,6,7 and 8) - but this needs confirming as calibration looks unreliable.

Effelsberg (A. Kraus at Fri Mar 11 12:51:39 2005) Minor failures : WIND
Snow (what else?) stopped us between UT2100-0000, 0026-0215, 0500-0510.